assuming that a couple are eqyually likely to produce a girls or a boy. find a probablility that in a family of 5 children there will be more boys than girls.
1
Expert's answer
2020-07-06T20:16:53-0400

Given : n = 5

P(boy)= 0.5

Since the experiment consists of n identical trials. Each trial results in one of the two outcomes, called success and failure. The probability of success, denoted p, remains the same from trial to trial, we use binomial distribution here.


To find : number of boys will be more when there are 3 and more than 3 boys.


That is P(X"\geq" 3)= 1-P(X<3)

= 1- P(x"\leq" 2)

Using binomial table ,

=1- 0.5

= 0.5